A charged particle  of charge 2×10-6 C and mass 10 milligram ,moving with a velocity of 1000m/s enters a uniform electric field of strength 103 NC-1 directed perpendicular to its direction of motion . Find the velocity and displacement of the particle after 10 s. 

Grade:12

1 Answers

Eshan
5 years ago
Dear student,

The component of velocity perpendicualr to the direction of electric field will remain constant since the particle does not accelerate in that direction.
Hencev_{perpendicular}=1000m/s \implies s_{perpendicular}=v_{perpendicular}t=10000m

The acceleration of particle in direction of electric field=\dfrac{qE}{m}=\dfrac{2\times 10^{-6}\times 10^3}{0.01}m/s^2=0.2m/s^2

Hence the velocity of particle parallel to electric field,v_{parallel}=at=2m/s
\implies s_{parallel}=\dfrac{1}{2}at^2=10m